A 2.86L sample of krypton gas contains a 0.403 mol of krypton. How many additional miles of krypton must be added to this sample to expand its volume to 19.9l assume temperature and pressure remain constant?
1
Expert's answer
2020-09-29T06:29:11-0400

From Avogadro's gas law, at constant temperature and pressure, V1/n1=V2/n2 Where V1= Initial Volume, n1= Initial amount/moles, V2= Final Volume, and n2= Final amount/moles.

From the data provided,

V1= 2.86 L

n1= 0.403 moles

V2= 19.9 L

n2= ?

From the law, n2= (V2x n1)/V1

=(19.9 x 0.403)/2.86

=2.804 moles.

Moles to be added = n2-n1

=2.804-0.403

= 2.401moles